A projectile leaves the ground at an angle 60 to the horizontal.its kinetic energy is E . neglecting air resistance,find in terms of E it's kinetic energy at the highest point of the motion


1
Expert's answer
2021-03-09T13:18:31-0500

Explanations & Calculations

  • At the heighest point the velocity of the projectile is only the horizontal component.
  • It should be known that the horizontal component of the velocity of a projectile remains unchaged thus it is "\small v\cos\theta=v \cos 60=\large \frac{v}{2}"
  • Therefore,kinetic energy it has at the highest point is
  • "\qquad\qquad
\begin{aligned}
\small E_{up}&=\small\frac{m(v\cos\theta)^2}{2}\\
&=\small \frac{mv^2}{8}\cdots(1)
\end{aligned}"


  • At the start its energy was,

"\qquad\qquad
\begin{aligned}
\small E&=\small\frac{mv^2}{2}\cdots(2)

\end{aligned}"

  • By (1)/(2),

"\qquad\qquad
\begin{aligned}
\small \frac{E_{up}}{E} &=\small\frac{1}{4}\\
\small E_{up}&=\small \frac{E}{4}
\end{aligned}"